OpenStudy (anonymous):

Sove for x: \[\log_2({8x})-\log_2({x^2-1})=\log_2{3}\]

OpenStudy (australopithecus):

combine the fractions and make both sides exponents to the base 2

OpenStudy (australopithecus):

i mean combine the logarithms

OpenStudy (australopithecus):

remember log(x) - log(y) is the same as log(x/y)

OpenStudy (australopithecus):

and also remember that \[2^{\log_2(x)} = x\]

OpenStudy (anonymous):

when you combine them do you add or subtract?

OpenStudy (australopithecus):

you don't add or subtract Look at this log(x) - log(y) is the same as log(x/y) Another example log(x-1) - log(y+2) is the same as log((x-1)/(y+2))

OpenStudy (australopithecus):

crud

OpenStudy (anonymous):

so it would look like \[\log_2(8x+x^2-1)=\log_23\]

OpenStudy (australopithecus):

here is another example of the second rule \[10^{\log(x)}\]

OpenStudy (australopithecus):

no

OpenStudy (australopithecus):

are you sure you are looking at the rule

OpenStudy (australopithecus):

to explain the rule \[\log((x+1)^{-1}) = -1\log(x+1) = - \log(x +1)\] Remembering that \[x^{-1} = 1/x\] thus you are simply putting the exponent back and multiplying the exponents in your case \[\log_2(8x)*\log_2((x^{2}-1)^{-1}) = \log_2(8x)*\log_2(1/(x^{2}-1)) = \log_2(8x/(x^{2}-1))\]

OpenStudy (australopithecus):

this is the mathematical reasoning behind the rule -log(x) = log(1/x) or log(x) - log(x) = log(x/x)

OpenStudy (australopithecus):

-log(22 + x)-log(2 + x) = log(1/(2+x)(22+x))

OpenStudy (australopithecus):

do you see it yet?

OpenStudy (anonymous):

so it would be 

OpenStudy (australopithecus):

\[\log_2(8x/(x^{2} - 1)) = \log_2(3) \] = \[2^{\log_2(8x/(x^{2} - 1))} = 2^{\log_2(3)}\] = 8x/(x^(2) - 1) = 3

OpenStudy (australopithecus):

no there is no negative in front of \[\log_2(8x)\]

OpenStudy (australopithecus):

so it is to the exponent +1 remember x^(+1) = x

OpenStudy (australopithecus):

http://www.sosmath.com/algebra/logs/log4/log44/log44.html here